OpenStudy (anonymous):

Simplify with math and/or identities. The equation is: cot(-B)sec(-B)sin(-B) So far I have -cot B sec B -sin B.

OpenStudy (anonymous):

cotB=(cosB/sinB) secB=(1/cosB) use these to arrive at your answer

OpenStudy (anonymous):

\[\cot(-B) = - \cot(B)\] \[\sec(-B) = \sec(B)\] \[\sin(-B) = - \sin(B)\] And then use : \[\cot(B) = \frac{\cos(B)}{\sin(B)}\] \[\sec(B) = \frac{1}{\cos(B)}\] After then, just simplify, cancel whatever is being cancelled...

OpenStudy (anonymous):

Thank you